OpenStudy (anonymous):

What is the standard form of 5x+5y=3

OpenStudy (abb0t):

What are you being asked to do?

OpenStudy (anonymous):

I'm being asked to find the slope and y-intercept

OpenStudy (anonymous):

@abb0t

OpenStudy (abb0t):

put it in the form \[y = mx+b\] in this case: m = slope b = y-value

OpenStudy (anonymous):

Slope-intercept form is y = mx + b, in that m is slope and b is y-intercept. All you have to do is start working towards putting everything on the right side, basically. 5x + 5y = 3 -5x -5x 5y = -5x + 3 -5 - 5 y = -5x - 2, if I'm not mistaken. With -2 as the y-intercept and -5 as the slope.

OpenStudy (abb0t):

@ClicheLlama I think it's best if you guide the person rather than do the work for them. That way they actually learn the material.